OpenStudy (moongazer):

the ave. of X numbers is 5. If Y is added to each number, the ave. will be? Please answer clearly and explain the solution clearly and easy to understand. Thanks! (guaranteed medal for a good answer)

OpenStudy (anonymous):

Not again...!

OpenStudy (moongazer):

you explained this already estudier but i forgot it

OpenStudy (moongazer):

i just have a question

OpenStudy (moongazer):

Hi moongazer, Average is sum/number, if you increase sum by 5y and divide by 5, then average is 5+y. (this is your answer before) my question is: could you make an equation as an explanation for this? not like the \[\sum_{?}^{?}\] equation)

OpenStudy (moongazer):

but if it easier to understand with that you may ^_^

OpenStudy (moongazer):

sorry for asking this question again :)

OpenStudy (anonymous):

Just do some algebra (a+b+c+d+e)/5 = 5 by definition of average. Then (a+b+c+d+e) = 25 (a+Y + b +Y +....) = 25 + 5Y so average is (25 +5Y)/5 = 5+Y

OpenStudy (anonymous):

U cannot make it any clearer than that...

OpenStudy (moongazer):

why you divided it by 5? (a+b+c+d+e)/5 = 5 by definition of average.

OpenStudy (anonymous):

by definition of average.

OpenStudy (anonymous):

The average is the sum divided by the number of items

OpenStudy (moongazer):

why is it it is 5? but the question says "X numbers" so the number of terms is X?

OpenStudy (anonymous):

I was using the answer I gave u as an example the last time to answer your question. Replace 5 with X and use...'s

OpenStudy (anonymous):

(a+b+c+d+e)/X = 5 by definition of average. Then (a+b+c+d+e) = 5X (a+Y + b +Y +....) = 5X + 5XY so average is (5X +5XY)/X = 5+Y

OpenStudy (anonymous):

Got it, now?

OpenStudy (anonymous):

Wait, mistake...

OpenStudy (anonymous):

(a+Y + b +Y +....) = 5X + XY so average is (5X +XY)/X = 5+Y

OpenStudy (anonymous):

I left the 5 in...